Magnetic moment of a bar magnet is M. the work done to turn the magnet by 90 degrees of magnet in direction of magnetic field B will be

Grade:12th pass

1 Answers

Eshan
5 years ago
Torque acting on a body of magnetic moment\vec{M}placed in a magnetic field\vec{B}is\vec{\tau}=\vec{M}\times \vec{B}=MB sin\theta.
Hence work done in turning bar magnetic from angle\theta_1to\theta_2is

W=\int_{\theta_1}^{\theta_2}MBsin\theta d\theta=MB(cos\theta_1-cos\theta_2)
Hence work done in turning bar magnet by\90^{\circ}=MB(cos0^{\circ}-cos90^{\circ})=MB